OpenStudy (anonymous):

Solve the following with a system of your choice 4x + 5y = 13 3x +y = -4

OpenStudy (anonymous):

solve the equations simultaneously

OpenStudy (anonymous):

What you mean?

OpenStudy (anonymous):

multiply 3x+y by 5

OpenStudy (anonymous):

5 (3x+y)

OpenStudy (anonymous):

then subtract it from the first equation

OpenStudy (anonymous):

sorry multiply 3x+y=-4 by 5

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

15x+5y=-20

OpenStudy (anonymous):

subtract 15x+5y=-20 from 4x+5y=13

OpenStudy (anonymous):

-11x=33

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

substitute the value of x in the equation

OpenStudy (anonymous):

-9+y=-4 y=5

OpenStudy (anonymous):

How did you get -9 + y = -4

OpenStudy (anonymous):

x=-3 substitute the value of x in the second equation we get -9+y=-4

OpenStudy (anonymous):

What do we do after that

OpenStudy (anonymous):

we get y=5

OpenStudy (anonymous):

y=-4+9 y=5

OpenStudy (anonymous):

Alright. can you help in in some other problems

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

fixin to post it